OpenStudy (kayders1997):

How do you take the derivative of g(x)=l(3x-4)l

OpenStudy (tkhunny):

Well, you don't for x = 4/3. For x > 4/3, the absolute values do nothing. Just discard them For x < 4/3, what? I'm not telling.

OpenStudy (kayders1997):

?

OpenStudy (tkhunny):

You'll have to do better than that. What part are you questioning?

OpenStudy (kayders1997):

what are you talking about?

OpenStudy (tkhunny):

What are you reading? It is a direct response to your question. Nothing tricky or evasive.

OpenStudy (kayders1997):

I don't understand what your trying to say I'm sorry I'm not smart at calculus

OpenStudy (doughnuttable):

Think about how the slope changes for l(3x-4)l as compared to 3x-4

OpenStudy (tkhunny):

What part of "You don't for x = 4/3." do you not understand?

OpenStudy (kayders1997):

it would always be positive, because when you take the absolute value the value can never be negitive

OpenStudy (doughnuttable):

Can you graph the slope?

OpenStudy (kayders1997):

Wouldn't the slope by 4/3

OpenStudy (tkhunny):

\(g(x) = |3x-4| = \begin{cases} 3x-4\;\;\;\;\;\;\;\text{if}\ x \ge 4/3 \\ -(3x-4)\;\text{if}\ x < 4/3 \end{cases}\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle |z|=\sqrt{z^2} }\) So, \(\color{#000000 }{ \displaystyle \frac{dy}{dx}|y|= \frac{d}{dx}\sqrt{y^2}=\frac{1}{2\sqrt{y^2}}\times{\color{blue}{2}y^{2-\color{red}{1}}} \times y' }\) \(\color{#000000 }{ \displaystyle \frac{dy}{dx}|y|= \frac{d}{dx}\sqrt{y^2}=\frac{2y\cdot y'}{2\sqrt{y^2}} =\frac{y}{|y|}\cdot y' }\)

OpenStudy (solomonzelman):

Just, by applying the definition of the absolute value (my first equation), I can conclude that, \(\color{#000000 }{ \displaystyle \frac{d}{dx}\left|f(x)\right| =\frac{f(x)\times f'(x)}{\left|f(x)\right|} }\)

OpenStudy (tkhunny):

In that case, you still don't for x = 4/3.

OpenStudy (doughnuttable):

You have to split the slope into two pieces because g(x) is not continuous.

OpenStudy (solomonzelman):

Yes, at the "vertex" (or whatever the corner is called), the function is not differentiable. try to draw the tangent line through that point (xD) :)

OpenStudy (kayders1997):

okay that makes sense, not that I think about it because at lxl you can't differentiate at (0,0)

OpenStudy (solomonzelman):

Yes, |x| you can't differentiate at x=0. You don't have an instantaneous slope there! However, You could get an explicit derivative-function for \(\small\color{#000000 }{ \displaystyle f(x)=\left|mx+b\right| }\), and note that for the derivative, you have a restriction, and that is: \(\color{#000000 }{ x\ne\frac{-b}{m} }\). But, for other than this, you just have, \(\color{#000000 }{ \displaystyle f'(x)=\frac{mx+b}{\left|mx+b\right|} \times m}\)

OpenStudy (kayders1997):

okay

OpenStudy (kayders1997):

is that right?